OpenStudy (anonymous):

5. Find the balance in the account to the nearest whole dollar: $2,500 principal earning 3%, compounded annually, after 4 years a. $2,575 b. $2,731 c. $2,814 d. $3,350

OpenStudy (sweetburger):

there is a equation usually for these types of questions do u happen to know it?

OpenStudy (anonymous):

A(4) = 2500(1+(0.03/4))^(3*4)? @sweetburger

OpenStudy (sweetburger):

A=P(1+r/n)^nt so if the interest is compounded annually its only compounded once a year meaning n=1 i believe

OpenStudy (sweetburger):

so A=2500(1+.03/1)^(1*4) I think this would get the complete correct answer

OpenStudy (anonymous):

Thank you! c: